# Uncovering and Mitigating Destructive Multi-Embedding Attacks in Deepfake Proactive Forensics

### v1

@@ Abstract
With the rapid evolution of deepfake technologies and the wide dissemination of digital media, personal privacy is facing increasingly serious security threats. Deepfake proactive forensics, which involves embedding imperceptible watermarks to enable reliable source tracking, serves as a crucial defense against these threats. Although existing methods show strong forensic ability, they rely on an idealized assumption of single watermark embedding, which proves impractical in real-world scenarios. In this paper, we formally define and demonstrate the existence of Multi-Embedding Attacks (MEA) for the first time. When a previously protected image undergoes additional rounds of watermark embedding, the original forensic watermark can be destroyed or removed, rendering the entire proactive forensic mechanism ineffective. To address this vulnerability, we propose a general training paradigm named Adversarial Interference Simulation (AIS). Rather than modifying the network architecture, AIS explicitly simulates MEA scenarios during fine-tuning and introduces a resilience-driven loss function to enforce the learning of sparse and stable watermark representations. Our method enables the model to maintain the ability to extract the original watermark correctly even after a second embedding. Extensive experiments demonstrate that our plug-and-play AIS training paradigm significantly enhances the robustness of various existing methods against MEA. Our code will be released upon acceptance.
1 Introduction
The rise of deepfake technology, particularly with the emergence of face manipulation models such as StarGAN [1] and LDM [2], has significantly lowered the barrier to content forgery, leading to severe societal risks. These technologies not only seriously erode the public’s trust in digital content, but also manipulate public opinion, infringe on personal privacy and encourage identity fraud. To address this growing challenge, digital forensics is shifting its focus from passive post-hoc detection [3, 4, 5] to proactive forensics [6, 7, 8], aiming to prevent malicious operations at their source. Currently, a leading strategy in deepfake proactive forensics involves embedding invisible digital watermarks [9] into original images, serving as a “digital fingerprint". When these images are disseminated or even subjected to deepfake manipulation, these embedded watermarks are designed to retain their integrity. This ultimately enables subsequent traceability and copyright protection. This concept is illustrated in Figure 1(a).
@@ 1 Introduction
Figure 1: Illustration of the ideal proactive forensics pipeline and the threat posed by Multi-Embedding Attacks (MEA). (a) In the standard pipeline, the embedded forensic watermark is expected to withstand manipulations such as deepfake generation. (b) However, additional watermark embeddings by third parties (e.g., social media platforms or malicious actors) can overwrite and degrade the original watermark, leading to a complete failure of the proactive forensics mechanism.
@@
Compared to traditional watermarking techniques [10, 11], deep watermarking methods based on encoder–decoder architectures have attracted increasing attention in proactive forensics due to their superior flexibility and embedding performance[12, 13]. Existing approaches have made significant progress in ensuring imperceptibility and robustness of watermarks against distortions [14]. However, existing approaches mainly focus on improving the robustness of watermarking schemes against various types of deepfake manipulations and common post-processing (e.g., adding noise), while ignoring the threat of multiple watermark embeddings. Just as facial images are vulnerable to deepfake forgery attacks, they are equally exposed to the risk of multiple watermark embedding operations during their widespread dissemination. Additional watermark embeddings can arise from automated processing by social media platforms. More significantly, malicious actors may deliberately introduce secondary watermarking to obscure the original forensic evidence. This phenomenon leaves a research blank to be solved urgently. That is, how can the forensic information of the original watermark be reliably extracted under the subsequent watermark embedding attack?
@@
To this end, we propose Adversarial Interference Simulation (AIS), a general and effective defense paradigm designed to against MEA. Rather than relying on complex architectural designs, AIS adjusts the optimization objective during the fine-tuning stage of the training process. At this stage, the training goal is explicitly defined as resisting interference from subsequent watermark embeddings, thereby preserving the original forensic information. Specifically, we explicitly apply simulated MEA to watermarked images and introduce a resilience loss to guide the encoder to learn more sparse watermark representation. Sparse watermarking has been widely validated as an effective strategy for enhancing robustness against attacks [15, 16]. Building on this foundation, we leverage sparse and stable watermark representations to resist MEA, ensuring the integrity of forensic information under adversarial interference. Crucially, as a model-agnostic training paradigm, AIS can be implemented as a plug-and-play solution to support a wide range of existing and future proactive forensic systems against MEA.
@@
2 Related Works
Existing passive deepfake detectors often struggle to generalize to unseen manipulation techniques[17, 18], prompting a growing interest in proactive forensic methods [19, 20]. Among these, invisible watermarking based on deep learning has become the dominant paradigm, typically employing an encoder–decoder architecture [21]. The robustness of such methods is commonly evaluated against deepfake attacks generated by techniques such as SimSwap [22], GANimation [23], and StarGAN [1]. A variety of sophisticated techniques have been developed to address these known threats, and significant advances have been made. For example, LampMark [7] introduces content-aware watermarking guided by facial landmarks to improve resistance. AdvMark [24] adopts adversarial watermarking strategies to improve both passive detection and active traceability under challenging conditions. Furthermore, recent work [25] introduces an unpredictable encryption mechanism to strengthen the security and confidentiality of watermark information. Although existing methods perform well in resisting content-level manipulations, they exhibit inherent limitations in both architectural design and optimization objectives. These deficiencies undermine their robustness against structured signal interference introduced by additional watermark embeddings.
@@ 2 Related Works
3 Analysis of Multi-Embedding Attacks
3.1 Problem Formulation
@@ 3.2 Theoretical Analysis of MEA
3.3 Empirical Validation of MEA
To validate our theoretical analysis, we provide empirical evidence supporting our claim that MEA constitutes a general vulnerability inherent in existing proactive forensic models, including SepMark [6], LampMark [7], WaveGuard [26], EditGuard [27]. In addition, due to the limited availability of open-source proactive forensic methods, we include two representative end-to-end robust watermarking methods, MBRS [12] and HiDDeN [28], in our evaluation.
@@ 4 Methodology
4.1 Adversarial Interference Simulation
4.1.1 Learning Sparse Representations.
@@ 4.1 Adversarial Interference Simulation
The standard task Loss ensures the model’s efficacy for a single watermark embedding. While specific components may vary slightly across different baseline architectures (e.g., SepMark), it consistently encompasses three fundamental objectives: (1) an encoder loss $\mathcal{L}_{I}$ for imperceptibility; (2) a multi-headed decoder loss $\mathcal{L}_{D}$ for basic robustness; (3) an adversarial loss $\mathcal{L}_{A}$ to further enhance stealthiness. We can thus define the standard task loss in its general form as:
@@
$\mathcal{L}_{Total}=\mathcal{L}_{task}+\lambda_{A}\mathcal{L}_{AIS}$ (8)
@@
5 Experiments
In this section, we conduct a series of comprehensive experiments to validate the effectiveness of our proposed AIS paradigm. More importantly, by highlighting the universal failure of existing methods against MEA and the success of our approach, we aim to establish MEA robustness as a necessary benchmark for future proactive forensic research.
5.1 Datasets and Implementation Details
5.1.1 Dataset and Baselines.
We conducted experiments on the widely used high-quality CelebA-HQ [29] dataset with a resolution of 256×256. To comprehensively evaluate the threat posed by MEA and the generalizability of our proposed AIS paradigm, we select a set of representative baseline methods, including SepMark [6], LampMark [7], WaveGuard [26], EditGuard [27], MBRS [12], and HiDDeN [28].
5.1.2 Implementation details.
All experiments were carried out using the PyTorch [30] framework on NVIDIA RTX 3090 GPU. The Adam optimizer [31] is used in all settings. During AIS fine-tuning, we follow the original hyperparameter settings of each baseline to ensure consistency. Detailed hyperparameter settings and training schedules for all baseline models can be found in appendix.
5.1.3 Evaluation metrics.
We assess performance using a comprehensive set of metrics. The primary indicator of robustness is the bit error rate (BER), which measures the integrity of the recovered forensic information; lower BER values indicate more successful recovery. To evaluate the visual quality of the watermarked images, we use the average peak signal-to-noise ratio (PSNR) and the structural similarity index measure (SSIM) [32]. These metrics quantify signal distortion and structural fidelity, respectively, with higher values reflecting less perceptible degradation caused by watermark embeddings.
5.2 Robustness to Intra-model MEA
To simulate scenarios in which adversaries deliberately disrupt forensic watermarks, we consider Intra-model MEA, a form of attack that poses significant threats to forensic methods. Detailed experimental procedures are provided in our appendix. As shown in Table 1, the experimental results clearly validate the effectiveness of the proposed AIS paradigm. Without AIS, the BER of baseline models increases sharply under repeated embeddings, approaching 50%, which reflects a complete breakdown of their forensic capabilities. In contrast, models enhanced with AIS consistently maintain low BERs even after five successive attacks, demonstrating strong resilience. This robustness is achieved without sacrificing visual quality, as indicated by consistently high PSNR and SSIM scores. These findings confirm that AIS can be seamlessly integrated into existing methods and serves as a plug-and-play enhancement that significantly improves robustness against MEA.
5.3 Generalization to Cross-model MEA
@@ 5.4 Ablation Study
5.5 Visual Quality Analysis
As corroborated by the high PSNR and SSIM values reported in Table 1, our AIS paradigm maintains excellent visual quality. Figure 4 provides a direct qualitative confirmation of this. We can clearly observe that the watermark images generated by all methods are visually indistinguishable from the original images, whether they are enhanced by AIS or not. This proves that our AIS paradigm significantly improves the security without sacrificing the most basic imperceptibility requirement of the watermark, which is completely consistent with the conclusion of objective indicators. Due to varying watermark embedding strategies, the magnified residuals exhibit distinct visual patterns. Additional examples are provided in the appendix.
6 Conclusion

### v2

@@ Abstract
With the rapid evolution of deepfake technologies and the wide dissemination of digital media, personal privacy is facing increasingly serious security threats. Deepfake proactive forensics, which involves embedding imperceptible watermarks to enable reliable source tracking, serves as a crucial defense against these threats. Although existing methods show strong forensic ability, they rely on an idealized assumption of single watermark embedding, which proves impractical in real-world scenarios. In this paper, we formally define and demonstrate the existence of Multi-Embedding Attacks (MEA) for the first time. When a previously protected image undergoes additional rounds of watermark embedding, the original forensic watermark can be destroyed or removed, rendering the entire proactive forensic mechanism ineffective. To address this vulnerability, we propose a general training paradigm named Adversarial Interference Simulation (AIS). Rather than modifying the network architecture, AIS explicitly simulates MEA scenarios during fine-tuning and introduces a resilience-driven loss function to enforce the learning of sparse and stable watermark representations. Our method enables the model to maintain the ability to extract the original watermark correctly even after a second embedding. Extensive experiments demonstrate that our plug-and-play AIS training paradigm significantly enhances the robustness of various existing methods against MEA.
Codehttps://github.com/vpsg-research/MEA
Introduction
@@
Figure 1: Illustration of the ideal proactive forensics pipeline and the threat posed by Multi-Embedding Attacks (MEA). (a) In the standard pipeline, the embedded forensic watermark is expected to withstand manipulations such as deepfake generation. (b) However, additional third-party embeddings (e.g., social media platforms or malicious actors) can overwrite and degrade the original watermark, leading to a complete failure of the proactive forensics mechanism.
The rise of deepfake technology, particularly with the emergence of face manipulation models such as StarGAN (choi2020stargan) and LDM (ldm), has significantly lowered the barrier to content forgery, leading to severe societal risks. These technologies not only seriously erode the public’s trust in digital content, but also manipulate public opinion, infringe on personal privacy and encourage identity fraud. To address this growing challenge, digital forensics is shifting its focus from passive post-hoc detection (Sun2021ltw; Exploit; Exposing) to proactive forensics (wu2023sepmark; wang2024lampmark; sun2025diffmark), aiming to prevent malicious operations at their source. Currently, a leading strategy in deepfake proactive forensics involves embedding invisible digital watermarks (chang2022blind) into original images, serving as a “digital fingerprint”. When these images are disseminated or even subjected to deepfake manipulation, these embedded watermarks are designed to retain their integrity. This ultimately enables subsequent traceability and copyright protection. This concept is illustrated in Figure 1(a).
@@
Compared to traditional watermarking techniques (sy2020traditionwatermarkcnn; wang2021mappingwatermarkcnn), deep watermarking methods based on encoder–decoder architectures have attracted increasing attention in proactive forensics due to their superior flexibility and embedding performance(jia2021mbrs; neekhara2022faceswatermarks). Existing approaches have made significant progress in ensuring imperceptibility and robustness of watermarks against distortions (wang2024robustmark). However, existing approaches mainly focus on improving the robustness of watermarking schemes against various types of deepfake manipulations and common post-processing (e.g., adding noise), while ignoring the threat of multiple watermark embeddings. Just as facial images are vulnerable to deepfake forgery attacks, they are equally exposed to the risk of multiple watermark embedding operations during their widespread dissemination. Additional watermark embeddings can arise from automated processing by social media platforms. More significantly, malicious actors may deliberately introduce secondary watermarking to obscure the original forensic evidence. This phenomenon leaves a research blank to be solved urgently. That is, how can the forensic information of the original watermark be reliably extracted under the subsequent watermark embedding attack?
@@
To this end, we propose Adversarial Interference Simulation (AIS), a general and effective defense paradigm designed to against MEA. Rather than relying on complex architectural designs, AIS adjusts the optimization objective during the fine-tuning stage of the training process. At this stage, the training goal is explicitly defined as resisting interference from subsequent watermark embeddings, thereby preserving the original forensic information. Specifically, we explicitly apply simulated MEA to watermarked images and introduce a resilience loss to guide the encoder to learn more sparse watermark representation. Sparse watermarking has been widely validated as an effective strategy for enhancing robustness against attacks (Boes2022Securesparse; Deeba2020sparsewatermark). Building on this foundation, we leverage sparse and stable watermark representations to resist MEA, ensuring the integrity of forensic information under adversarial interference. Crucially, as a model-agnostic training paradigm, AIS can be implemented as a plug-and-play solution to support a wide range of existing and future proactive forensic systems against MEA.
@@
Related Works
Existing passive deepfake detectors often struggle to generalize to unseen manipulation techniques(liu2024faceforgerydetection; xu2024exposingdeepfake; guo2023constructing; LDFnet), prompting a growing interest in proactive forensic methods (yu2021fingerprinting; yu2022responsible; he2025kad). Among these, invisible watermarking based on deep learning has become the dominant paradigm, typically employing an encoder–decoder architecture (yang2021faceguard). The robustness of such methods is commonly evaluated against deepfake attacks generated by techniques such as SimSwap (chen2020simswap), GANimation (gao2021information), and StarGAN (choi2020stargan). A variety of sophisticated techniques have been developed to address these known threats, and significant advances have been made. For example, LampMark (wang2024lampmark) introduces content-aware watermarking guided by facial landmarks to improve resistance. AdvMark (Wu2024Advmark) adopts adversarial watermarking strategies to improve both passive detection and active traceability under challenging conditions. Furthermore, recent work (wang2024robustmark) introduces an unpredictable encryption mechanism to strengthen the security and confidentiality of watermark information. Although existing methods perform well in resisting content-level manipulations, they exhibit inherent limitations in both architectural design and optimization objectives. These deficiencies undermine their robustness against structured signal interference introduced by additional watermark embeddings.
@@
Analysis of Multi-Embedding Attacks
Problem Formulation
@@
Empirical Validation of MEA
To validate our theoretical analysis, we provide empirical evidence supporting our claim that MEA constitutes a general vulnerability inherent in existing proactive forensic models, including SepMark (wu2023sepmark), LampMark (wang2024lampmark), WaveGuard (he2025waveguard), EditGuard (zhang2024editguard). In addition, due to the limited availability of open-source proactive forensic methods, we include two representative end-to-end robust watermarking methods, MBRS (jia2021mbrs) and HiDDeN (zhu2018hidden), in our evaluation.
@@
Adversarial Interference Simulation
Learning Sparse Representations.
@@
The standard task Loss ensures the model’s efficacy for a single watermark embedding. While specific components may vary slightly across different baseline architectures (e.g., SepMark), it consistently encompasses three fundamental objectives: (1) an encoder loss $\mathcal{L}_{I}$ for imperceptibility; (2) a multi-headed decoder loss $\mathcal{L}_{D}$ for basic robustness; (3) an adversarial loss $\mathcal{L}_{A}$ to further enhance stealthiness. The general form of the task loss is therefore defined as:
@@
$\mathcal{L}_{Total}=\mathcal{L}_{task}+\lambda_{A}\mathcal{L}_{AIS}.$ (8)
@@
Experiments
Datasets and Implementation Details
Dataset and Baselines.
We conducted experiments on the the widely used CelebA-HQ (karras2018celebhq) dataset (256×256), and further performed cross-dataset evaluation on the LFW (huang2008lfw) dataset. To comprehensively evaluate the threat posed by MEA and the generalizability of our proposed AIS paradigm, we select a set of representative baseline methods, including SepMark (wu2023sepmark), LampMark (wang2024lampmark), WaveGuard (he2025waveguard), EditGuard (zhang2024editguard), MBRS (jia2021mbrs), and HiDDeN (zhu2018hidden).
Implementation details.
All experiments were carried out using the PyTorch (paszke2019pytorch) framework on NVIDIA RTX 3090. The Adam optimizer (kingma2014adam) is used in all settings. During AIS fine-tuning, we follow the original hyperparameter settings of each baseline to ensure consistency. Detailed hyperparameter settings and training schedules for all models can be found in appendix.
Evaluation metrics.
We assess performance using a comprehensive set of metrics. The primary indicator of robustness is the bit error rate (BER), which measures the integrity of the recovered forensic information; lower BER values indicate more successful recovery. To evaluate the visual quality of the watermarked images, we use the average peak signal-to-noise ratio (PSNR) and the structural similarity index measure (SSIM) (wang2004ssim). These metrics quantify signal distortion and structural fidelity, respectively, with higher values reflecting less perceptible degradation caused by watermark embeddings.
Robustness to Intra-model MEA
To simulate scenarios in which adversaries deliberately disrupt forensic watermarks, we consider Intra-model MEA, a form of attack that poses significant threats to forensic methods. Table 1 validates the effectiveness the proposed AIS paradigm. Without AIS, the BER of baseline models rises sharply under repeated embeddings, approaching 50% and indicating a near-complete failure of forensic functionality. In contrast, models enhanced with AIS consistently maintain low BERs even after five successive attacks, showing strong resilience. This robustness is achieved without sacrificing visual quality, as indicated by consistently high PSNR and SSIM scores. These results indicate that AIS can be seamlessly integrated into existing methods as a plug-and-play module that markedly improves robustness against MEA. To examine cross-dataset generalization, we also evaluate the LTW dataset, with results presented in Table 4.
Metric Emb. SepMark EditGuard MBRS
BER (%) $\downarrow$
1 0.49 / 3.17 0.02 / 8.29 0.20 / 0.58
2 49.27 / 20.16 43.41 / 12.93 48.26 / 1.05
PSNR $\uparrow$
1 25.50 / 25.98 42.57 / 48.12 38.06 / 38.01
2 13.29 / 11.29 38.62 / 43.76 34.73 / 34.38
SSIM $\uparrow$
1 0.69/ 0.71 0.99 / 1.00 0.93 / 0.92
2 0.22 / 0.15 0.98 / 0.99 0.88 / 0.88
Table 4: Performance comparison on the LFW dataset. Results are reported in the format (Baseline / +AIS).
Generalization to Cross-model MEA
@@
Visual Quality Analysis
AIS maintains excellent visual quality, confirmed by high PSNR and SSIM scores in Table 1. Figure 4 provides a direct qualitative confirmation. We can clearly observe that the watermark images generated by all methods are visually indistinguishable from the original images, whether they are enhanced by AIS or not. This proves that our AIS paradigm significantly improves the security without sacrificing the most basic imperceptibility requirement of the watermark, which is completely consistent with the conclusion of objective indicators. Due to varying watermark embedding strategies, the magnified residuals exhibit distinct visual patterns.
Conclusion
@@
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant 62462060, Grant 62302427, and Grant 62472368, in part by the Natural Science Foundation of Xinjiang Uygur Autonomous Region under Grant 2023D01C175.